### v1

@@
(April 4, 2024)
This paper presents the first quasi-experimental research examining the effect of both local and state anti-discrimination laws on sexual orientation on the labor supply and wages of lesbian, gay, and bisexual (LGB) workers. To do so, we use the American Community Survey data on household composition to infer sexual orientation and combine this with a unique panel dataset on local anti-discrimination laws. Using variation in law implementation across localities over time and between same-sex and different-sex couples, we find that anti-discrimination laws significantly reduce gaps in labor force participation rate, employment, and the wage gap for gay men relative to straight men. These laws also significantly reduce the labor force participation rate, employment, and wage premium for lesbian women relative to straight women. One explanation for the reduced labor supply and wage premium is that lesbian couples begin to have more children in response to the laws. Finally, we present evidence that state anti-discrimination laws significantly and persistently increased support for same-sex marriage. This research shows that anti-discrimination laws can be an effective policy tool for reducing labor market inequalities across sexual orientation and improving sentiment toward LGB Americans.
@@ 1 Introduction
On April 15, 2020, the Supreme Court of the United States ruled that federal employment discrimination law applied to lesbian, gay, bisexual, and transgender (LGBT) individuals. Before this ruling, the legislation in 25 states made it legal for private businesses to fire an employee for being LGBT. Unlike race, sex, age, and religion, federal law did not include sexual orientation as a protected class for federal anti-discrimination laws until 2020. As a result, laws extending protected class status based on sexual orientation varied between states and localities. The first state to pass an anti-discrimination law based on sexual orientation was Wisconsin in 1982, and the most recent state to pass such a law was Pennsylvania and Michigan in 2023.111Since August 2018, Pennsylvania has interpreted discrimination based on sexual orientation and gender identity as being covered under the “sex” category of the Pennsylvania Human Relations Act. This allowed LGBT individuals to file complaints for discrimination in employment, housing, education, and public accommodations. Following Michigan’s lead in May 2018, Pennsylvania became the second state to implement statewide LGBT protections in this manner. There has been a renewed push to enact sexual orientation protections at the legislative level as Democrats in the House of Representatives introduced “The Equality Act” in early 2019, which would add sexual orientation and gender identity as federally protected characteristics. The focus of this paper is the effect of local and state sexual orientation anti-discrimination legislation on labor market differentials for lesbian, gay, and bisexual (LGB) workers.
We exploit the differential roll-out of state and local laws from 2005-2019 in a modern triple difference-in-differences framework to analyze how these anti-discrimination laws impact the wages and labor supply of LGB workers. The economics literature on LGB workers consistently finds that gay/bisexual men have a pay and labor supply penalty and lesbian/bisexual women have a pay and labor supply premium over their heterosexual counterparts222(Badgett (1995); Klawitter and Flatt (1998); Allegretto and Arthur (2001); Black et al. (2003); Carpenter (2005); Black et al. (2007); Antecol et al. (2008); Klawitter (2015); Jepsen and Jepsen (2017); Carpenter and Eppink (2017)). The differences in pay are attributed to a host of factors including discrimination and intra-household labor allocation and specialization. We test the role of discrimination by examining how the passage of anti-discrimination laws affects the labor supply and wage differences between LGB and heterosexual workers.
@@
Previous research has leveraged state-level sexual orientation anti-discrimination laws within a difference-in-differences framework. [^4] The Williams Institute found that sexual orientation discrimination occurs at a similar rate to sex-based discrimination using state-level complaint data following state anti-discrimination laws (Ramos et al., 2008). Martell (2013) study demonstrated that such laws have significantly reduced wage disparities for gay men by 20%.555 Since Martell (2013) and Klawitter and Flatt (1998) there has been a booming literature on difference-in-differences. Accordingly, unlike Martell (2013) and Klawitter and Flatt (1998), we employ modern difference-in-difference estimators to alleviate biases stemming from staggered policy design. Our research extends beyond the scope of previous studies by incorporating an examination of local laws, thereby offering a more comprehensive understanding of the legislative landscape. Furthermore, this study also estimates the effects of such laws on women, a demographic previously under-explored in this context. Our study expands the scope of Klawitter and Flatt (1998) by leveraging a larger dataset, utilizing American Community Survey (ACS) data spanning from 2004 to 2019, in contrast to Klawitter and Flatt (1998) which relied on the 1990 U.S. census data encompassing approximately 13,000 married couples, 14,500 different-sex unmarried couples, and 6,800 same-sex couples. The ACS data allows for a robust analysis across a broader temporal spectrum, capturing a wide array of demographic and economic changes over 15 years. This extensive dataset, combined with the application of modern difference-in-difference estimators, enables a more detailed and dynamic examination of the effects of anti-discrimination policies on the earnings and household incomes of same-sex couples. Additionally, Burn (2018)’s more recent examination in a triple difference-in-differences framework confirms these findings but also underscores a gap in the literature: the lack of consideration for local anti-discrimination ordinances. Previous studies have not fully addressed potential biases arising from variations in the timing of these laws’ enactment. To bridge this gap, our research employs the methodology proposed by Borusyak et al. (2021), comparing it with the conventional triple difference-in-differences estimator, finding both approaches yield similar results.
@@
We use the 2005-2019 American Community Survey (ACS) and household composition to infer sexual orientation and create a unique and novel panel dataset on the passage of local and state anti-discrimination laws. We collected information on local laws from a host of sources including media reports, FOIA requests, and an advocacy group. We find a significant reduction in differences between LGB workers and heterosexual workers across labor supply and wage measures, due to anti-discrimination laws. Anti-discrimination laws significantly reduce the gap in labor force participation and employment of gay men by 2.1 p.p. and 1.5 p.p., respectively. The laws also significantly increase hourly wages for gay men by 6.2%. The results differ for women, with the laws significantly reducing their labor force participation and employment premium over straight women by 1.7 p.p., and 2.3 p.p., respectively. Using an event study plot, we show that the outcome trends are parallel before the implementation of the anti-discrimination laws suggesting the workers in treatment and control regions are reasonable comparisons. [^7] In Appendix A, we replicate the findings of past research that finds gay/bisexual men have lower labor force participation and employment rates and make 8-11% less than their employed straight counterparts using hourly wages and annual earnings. We also replicate findings that lesbian/bisexual women have a higher labor force participation and employment rate and earn 5-15% more than their employed straight counterparts.
We explore theories for the differing effects of anti-discrimination laws on gay men and lesbian women in the Discussion section, using the Becker (1981) model of household specialization. In the traditional Becker (1981) model of household specialization, men typically specialize in market production, and women typically specialize in household production, in part due to differences in biology where women birth and care for children, resulting in a one-earner household. These differences in household specialization are less pronounced in same-sex partnerships, but they may become more similar for women in same-sex partnerships following the passage of an anti-discrimination law if it gives greater protection to the higher wage earner. We show empirically that the difference in hours worked between partners within a lesbian household goes up relative to gay households following anti-discrimination laws, suggesting that lesbian households could become more specialized following anti-discrimination laws with one woman working more hours and the other woman working fewer to focus on household production. We also show that lesbian households have significantly more children than gay households after the passage of an anti-discrimination law. More children could induce lesbian households to further specialize in the intra-household division of labor, adopting a more traditional household model to help care for additional children.
@@ 2 Data
A common issue in analyzing the pay gap/premium between homosexual and heterosexual workers is a lack of high-quality data about sexual orientation, wages, and employment. We follow the literature (Klawitter and Flatt (1998); Gates (2009); Jepsen and Jepsen (2017)) in inferring sexual orientation by looking at household composition. Specifically, we infer a person’s sexual orientation to be homosexual/bisexual if they have an “unmarried partner” or a spouse that is the same sex as themselves. [^8] We exclude individuals whose sex or relationship status was allocated following (Gates and Steinberger, 2009).. The ACS references a close personal relationship for unmarried partners as well as giving an option for “partner/roommate”, which reduces the likelihood that straight roommates would misclassify as unmarried partners.
@@
The comparisons that we make are between people in same-sex relationships, defined as being in an unmarried partnership or married with someone of the same sex, and those in different-sex relationships. This comparison based on household composition, though standard in the literature, is not equivalent to comparing LGB and heterosexual workers. It compares those who are in a same-sex relationship to those who are in a different-sex relationship. Notably, this comparison excludes all workers who are not in a cohabitating relationship and classifies bisexual people in different-sex relationships as members of the control group. For example, if a person is misclassified as straight instead of bisexual, and bisexual workers have lower pay than their straight counterparts, then the misclassification would bias our pay gap/premium results toward zero.
@@
We use the 2005-2019 yearly ACS from (Ruggles et al., 2019). We use wages, defined as salaried wages from an employer. The ACS reports wage earnings in annual terms. We use the annual figures and convert them to hourly wage figures using variables on average weekly hours worked and weeks worked999“Weeks worked” in the ACS is a categorical variable giving a range of weeks worked. We take the median value given in the range to compute weeks worked.. Finally, we limit our sample to prime-age working adults and only examine those who are ages 25-65. We use 25 as the lower age cutoff to allow workers to complete college and have more fully developed human capital, which is typically thought of as an important factor for the differences in wages between LGB and straight workers (Black et al., 2007).
@@
We obtained data on the passage of state anti-discrimination laws from LGBTMap.org, an LGBT advocacy group. The website gives information on which states passed anti-discrimination laws and when. We focus solely on sexual orientation anti-discrimination laws that give protection in employment. LGBTMap.org also provided incomplete data detailing the passage of local city and county laws with many cities missing years for the start of their anti-discrimination laws. We completed the dataset for the relevant years using old media reports, correspondence with local officials, and FOIA requests. This dataset is the first complete city-level dataset on sexual orientation anti-discrimination laws in the U.S. We focus on cities reported in the ACS and match those city laws with their corresponding counties to merge with the ACS. The ACS only reports county of residence for those in metro areas, so any rural counties are lumped together. We make the assumption in our analysis that the impact of these laws is the same across states and localities, estimating an average effect.
@@
The enactment of anti-discrimination legislation is not random. The states that have these protections are generally considered friendlier to LGB workers than those without these laws and have a larger share of same-sex partnerships than those states without these laws. Also, many counties with anti-discrimination laws have large cities with a larger concentration of LGB workers than rural counties. However, this is not universal. For instance, Utah extended protection to LGB workers despite being a relatively conservative state, and certain liberal cities like Houston [^10] Houston passed a sexual orientation and gender identity anti-discrimination ordinance, but it was only in effect for 3 months before being challenged. The law was put up to a public vote and lost, repealing the law. are noticeably absent from extending protection to LGB workers. One potential concern in this analysis is the endogenous adoption of laws. The areas that are friendly and less hostile to LGB workers may be the areas that are most likely to adopt sexual orientation anti-discrimination laws. The main threat to identification is that the timing of anti-discrimination laws is correlated with another factor causing an omitted variable bias. It could be that positive sentiment toward LGB workers causes both the law change and any change in labor market outcomes. We mitigate this by adopting a saturated triple-difference model, including all two-way fixed effects (county-year, year-SSP, and county-SSP), as proposed by (Burn, 2018). The inclusion of county-year fixed effects allows us not to include county controls hence eliminating the need to control for unavailable local levels of support for same-sex marriage. We refine the approach of (Burn, 2018) by employing the methodology of (Borusyak et al., 2021), particularly suited for analyzing staggered policy adoptions like ours. This method accurately estimates the effects of anti-discrimination laws, considering their varied implementation times. For additional validation, we include results from a simpler difference-in-differences estimator in the appendix, providing robustness to our findings.
@@
In Figure 2, we present the distribution of same-sex partnerships by state from 2005 to 2019, and in Table 2, we present the 10 counties with the largest share of same-sex partnerships. Unsurprisingly, the distribution of same-sex partnerships by state and by county is skewed toward more progressive states and counties with large cities in them that are known for having a large LGB population like San Francisco, the District of Columbia, New York City, and Boston. The LGB population varies from the straight population in many dimensions including geography.
@@
We present descriptive statistics showing the differences in the LGB population compared to the straight population broken down by education level and sex in Table 3. Descriptively, there is a clear difference in labor market outcomes and characteristics between men and women in same-sex partnerships and those in different-sex partnerships when controlling for education.
@@ 3 Identification Strategy & Estimation
$\begin{split}Y_{ict}=\alpha+\beta_{0}(SSP_{ict}\times ENDA_{ct}\times I_{\textrm{Sex=Female})}+\beta_{1}(SSP_{ict}\times ENDA_{ct}\times I_{\textrm{Sex=Male})}\\ +(I_{c}\times I_{t}\times I_{\textrm{Sex=Male}}))\gamma_{sct}+(SSP_{ict}\times I_{c}\times I_{\textrm{Sex=Male}}))\theta_{S,s,c}\\ +(SSP_{ict}\times I_{t}I_{\textrm{Sex=Male}}))\theta_{S,s,t}+X^{\prime}_{ict}\delta+\epsilon_{ict}\end{split}$ (1)
@@
In our regression, equation (1) defines the labor market outcomes $Y_{ict}$ for individual $i$, in county $c$, at time $t$ with sex $s$ and same-sex partnership status $S$. These include the labor force participation, employment status, or the real value of wages, in 1999 dollars. The binary variable $SSP_{ict}$ indicates if an individual is in a same-sex partnership. The model explores the interaction between being in a same-sex partnership and the presence of county-level Employment Non-Discrimination Acts $(SSP_{ict}\times ENDA_{ct})$. [^11] Treatment status is computed using the earliest year of ENDAs at the county or state level, with individuals remaining classified as “treated” thereafter throughout the dataset. Given the ACS’s 23-month reference period for earnings, as discussed by Hansen et al. (2020), there’s a potential for bias in treatment assignment. This extended timeframe might not precisely align with the exact implementation dates of ENDAs, potentially biasing the immediate effects of these policies towards zero. The fixed effects for county-year-sex interactions $(\gamma_{ct})$, control for unobserved heterogeneity. Interactions between the same-sex partnership status, sex and county $(SSP_{ict}\times I_{c}\times I_{\textrm{Sex=Male}})$ and year fixed effects $(SSP_{ict}\times I_{t}\times I_{\textrm{Sex=Male}})$ capture varying impacts over time and across different regions by sex. $X^{\prime}_{ict}$ is a vector of person-specific covariates including race, children, age, and education level - each represented up to a quadratic term to capture potential nonlinear effects. The error term $\epsilon_{ict}$ is clustered at the county level to account for within-group correlation. [^12] Since the ACS only identifies metro counties, the county fixed effect is close to a pseudo-county fixed effect. The rural counties are all compared together with a true county fixed effect for metro area counties. This specification aims to isolate the effects of anti-discrimination laws on the labor market outcomes for individuals in same-sex partnerships by sex.
Our sample is limited to people aged 25 to 65, and we estimate this model jointly by sex using the ACS person weights. When examining wages, the $\ln(Y_{ict}+1)$ specification is adopted to include both employed and unemployed individuals in the analysis, recognizing employment as endogenous to anti-discrimination laws. This approach ensures a comprehensive evaluation of the law’s impact on the entire income spectrum.
In equation (1), the differential effect of being a cohabiting same-sex couple by sex in the United States over time is captured by $\theta_{S,s,t}$. The county fixed effects for cohabiting same-sex couples by sex $\theta_{S,s,c}$ capture how the wage differential varies across counties by sex. The parameters of interest is $\beta_{0}$ and $\beta_{1}$, which capture how the differential between homosexuals and heterosexuals in counties with ENDAs changed after the law’s passage by sex. Hence, $\beta_{1}$, will give the effect of ENDAs on the labor supply and pay gap/premium for males, and $\beta_{0}$ for females. [^13] We recognize the heterogeneous impact of anti-discrimination laws across different regions, yet our analysis aims to estimate an average effect for those directly subject to the laws.

### v2

@@
(April 1, 2025)
This paper presents the first quasi-experimental research examining the effect of both local and state anti-discrimination laws on sexual orientation on the labor supply and wages of lesbian, gay, and bisexual (LGB) workers. To do so, we use the American Community Survey data on household composition to infer sexual orientation and combine this with a unique panel dataset on state and local anti-discrimination laws. Leveraging variation in law implementation across localities over time and between same-sex and different-sex couples, we find that anti-discrimination laws not only significantly narrow gaps in labor force participation and employment for male same-sex couples relative to men in different-sex couples, but also boost their percentile rank in the wage distribution. Our analysis reveals mostly null effects for female same-sex couples; however, in metropolitan areas these laws significantly reduce their employment compared to women in different-sex couples. One explanation for the reduced labor supply is that female same-sex couples begin to have more children in response to the laws. Finally, we present evidence that state anti-discrimination laws significantly and persistently increased support for same-sex marriage. This research shows that anti-discrimination laws can be an effective policy tool for reducing labor market inequalities across sexual orientation and improving sentiment toward LGB Americans.
@@ 1 Introduction
On April 15, 2020, the Supreme Court of the United States ruled that federal employment discrimination law applied to lesbian, gay, bisexual, and transgender (LGBT) individuals. Before this ruling, the legislation in 25 states made it legal for private businesses to fire an employee for being LGBT. Unlike race, sex, age, and religion, federal law did not include sexual orientation as a protected class for federal anti-discrimination laws until 2020. As a result, laws extending protected class status based on sexual orientation varied between states and localities. The first state to pass an anti-discrimination law based on sexual orientation was Wisconsin in 1982, and the most recent state to pass such a law was Pennsylvania and Michigan in 2023.111Since August 2018, Pennsylvania has interpreted discrimination based on sexual orientation and gender identity as being covered under the “sex” category of the Pennsylvania Human Relations Act. This allowed LGBT individuals to file complaints for discrimination in employment, housing, education, and public accommodations. Following Michigan’s lead in May 2018, Pennsylvania became the second state to implement statewide LGBT protections in this manner. There has been a renewed push to enact sexual orientation protections at the legislative level as Democrats in the House of Representatives introduced “The Equality Act” in early 2019, which would add sexual orientation and gender identity as federally protected characteristics. The focus of this paper is the effect of local and state sexual orientation anti-discrimination legislation on labor market differentials for lesbian, gay, and bisexual (LGB) workers. [^2] For a review of the impacts of legal access to same-sex marriage, see Badgett et al. (2024). For a broader review of the economics related to sexual orientation and gender identity, see Badgett et al. (2023).
We exploit the differential roll-out of state and local laws from 2005-2019 in a modern triple difference-in-differences framework to analyze how these anti-discrimination laws impact the wages and labor supply of LGB workers. The economics literature on LGB workers consistently finds that gay/bisexual men have a pay and labor supply penalty and lesbian/bisexual women have a pay and labor supply premium over their heterosexual counterparts333(Badgett (1995); Klawitter (2011); Klawitter and Flatt (1998); Allegretto and Arthur (2001); Black et al. (2003); Carpenter (2005); Black et al. (2007); Antecol et al. (2008); Klawitter (2015); Jepsen and Jepsen (2017); Carpenter and Eppink (2017)), Tilcsik et al. (2015). The differences in pay are attributed to a host of factors including discrimination and intra-household labor allocation and specialization. We test the role of discrimination by examining how the passage of anti-discrimination laws affects the labor supply and wage differences between LGB and heterosexual workers.
@@
Previous research has leveraged state-level sexual orientation anti-discrimination laws within a difference-in-differences framework. [^5] The Williams Institute found that sexual orientation discrimination occurs at a similar rate to sex-based discrimination using state-level complaint data following state anti-discrimination laws (Ramos et al., 2008). Martell (2013) study demonstrated that such laws have significantly reduced wage disparities for gay men by 20%.666 Since Martell (2013) and Klawitter and Flatt (1998) there has been a booming literature on difference-in-differences. Accordingly, unlike Martell (2013) and Klawitter and Flatt (1998), we employ modern difference-in-difference estimators to alleviate biases stemming from staggered policy design. Our research extends beyond the scope of previous studies by incorporating an examination of local laws, thereby offering a more comprehensive understanding of the legislative landscape. Furthermore, this study also estimates the effects of such laws on women, a demographic previously under-explored in this context. Our study expands the scope of Klawitter and Flatt (1998) by leveraging a larger dataset, utilizing American Community Survey (ACS) data spanning from 2005 to 2019, in contrast to Klawitter and Flatt (1998) which relied on the 1990 U.S. census data encompassing approximately 13,000 married couples, 14,500 different-sex unmarried couples, and 6,800 same-sex couples. The ACS data allows for a robust analysis across a broader temporal spectrum, capturing a wide array of demographic and economic changes over 15 years. This extensive dataset, combined with the application of modern difference-in-difference estimators, enables a more detailed and dynamic examination of the effects of anti-discrimination policies on the earnings and household incomes of same-sex couples. Additionally, Burn (2018)’s more recent examination in a triple difference-in-differences framework confirms these findings but also underscores a gap in the literature: the lack of consideration for local anti-discrimination ordinances. Previous studies have not fully addressed potential biases arising from variations in the timing of these laws’ enactment.
@@
We use the 2005-2019 American Community Survey (ACS) and household composition to infer sexual orientation and create a unique and novel panel dataset on the passage of local and state anti-discrimination laws. We collected information on local laws from a host of sources including media reports, FOIA requests, and an advocacy group. We find a significant reduction in differences between LGB workers and heterosexual workers across labor supply and wage measures, due to anti-discrimination laws. Anti-discrimination laws significantly reduce the gap in labor force participation and employment of men in male same-sex couples by 3.8 percentage points (78%) and 3.7 percentage points (77%), respectively. The laws improve men in male same-sex couples relative position in the wage distribution, with their percentile rank in the hourly wage and annual earnings distributions improving on average by approximately 2.7 and 3 percentiles, respectively. By contrast, the wage distribution for women in same-sex partnerships remains largely unaffected by these laws. Using an event study plot, we show that the outcome trends are parallel before the implementation of the anti-discrimination laws suggesting the workers in treatment and control regions are reasonable comparisons. [^8] In Appendix B A, we replicate the findings of past research that finds men in male same-sex couples have lower labor force participation and employment rates and make 8-11% less than their employed straight counterparts using hourly wages and annual earnings. We also replicate findings that women in female same-sex couples have a higher labor force participation and employment rate and earn 5-15% more than their employed straight counterparts.
We explore theories for the differing effects of anti-discrimination laws on men in male same-sex couples and women in female same-sex couples in the Discussion section. Reduced discrimination improves job security for sexual minorities (Badgett, 2006), which may allow the secondary earner to reduce market participation in favor of increased household production, including child rearing, a pattern consistent with household bargaining models (Chiappori, 1992; Lundberg and Pollak, 1993). These differences in household specialization are less pronounced in same-sex partnerships, but they may become more similar for women in female same-sex couples following the passage of an anti-discrimination law if it gives greater protection to the higher wage earner. We show empirically that female same-sex households have significantly more children than male same-sex households after the passage of an anti-discrimination law.
@@ 2 Data
A common issue in analyzing the pay disparities between homosexual and heterosexual workers is a lack of high-quality data about sexual orientation, wages, and employment. We follow the literature (Klawitter and Flatt (1998); Gates (2009); Jepsen and Jepsen (2017)) in inferring sexual orientation by looking at household composition. Specifically, we infer a person’s sexual orientation to be homosexual/bisexual if they have an “unmarried partner” or a spouse that is the same sex as themselves. [^9] We exclude individuals whose sex or relationship status was allocated following (Gates and Steinberger, 2009).. The ACS references a close personal relationship for unmarried partners as well as giving an option for “partner/roommate”, which reduces the likelihood that different-sex roommates would misclassify as unmarried partners.
@@
The comparisons that we make are between people in same-sex partnerships (SSP), defined as being in an unmarried partnership or married with someone of the same sex, and those in different-sex partnerships (DSP). This comparison based on household composition, though standard in the literature, is not equivalent to comparing LGB and heterosexual workers. It compares those who are in a same-sex relationship to those who are in a different-sex relationship. Importantly, this comparison excludes all workers who are not in a cohabitating relationship and classifies bisexual individuals in different-sex relationships as members of the control group. Because bisexual workers consistently earn less than both heterosexual and gay or lesbian counterparts ((Mize, 2016)), this classification introduces a bias toward underestimating pay disparities. Additionally, similar misclassification may occur if gay or lesbian individuals are categorized as heterosexual. Thus, the estimated pay gaps presented here likely represent conservative estimates.
@@
We use the 2005-2019 yearly ACS from (Ruggles et al., 2019). We use wages, defined as salaried wages from an employer. The ACS reports wage earnings in annual terms. We use the annual figures and convert them to hourly wage figures using variables on average weekly hours worked and weeks worked101010“Weeks worked” in the ACS is a categorical variable giving a range of weeks worked. We take the median value given in the range to compute weeks worked.. Finally, we limit our sample to prime-age working adults and only examine those who are ages 25-65. We use 25 as the lower age cutoff to allow workers to complete college and have more fully developed human capital, which is typically thought of as an important factor for the differences in wages between LGB and straight workers (Black et al., 2007).
@@
We obtained data on the passage of state anti-discrimination laws from LGBTMap.org, an LGBT advocacy group. The website gives information on which states passed anti-discrimination laws and when. We focus solely on sexual orientation anti-discrimination laws that give protection in employment. LGBTMap.org also provided incomplete data detailing the passage of local city and county laws with many cities missing years for the start of their anti-discrimination laws. We completed the dataset for the relevant years using old media reports, correspondence with local officials, and FOIA requests. Building on earlier city-level analyses such as (Klawitter, 2011), this dataset provides the first complete city-level dataset covering 15 years of sexual orientation anti-discrimination laws in the U.S. We focus on cities reported in the ACS and match those city laws with their corresponding counties to merge with the ACS. The ACS only reports county of residence for those in metro areas, so any rural counties are lumped together within a state.
@@
The enactment of anti-discrimination legislation is not random. The states that have these protections are generally considered friendlier to LGB workers than those without these laws and have a larger share of same-sex partnerships than those states without these laws. Also, many counties with anti-discrimination laws have large cities with a larger concentration of LGB workers than rural counties. However, this is not universal. For instance, Utah extended protection to LGB workers despite being a relatively conservative state, and certain liberal cities like Houston [^11] Houston passed a sexual orientation and gender identity anti-discrimination ordinance, but it was only in effect for 3 months before being challenged. The law was put up to a public vote and lost, repealing the law. are noticeably absent from extending protection to LGB workers. One potential concern in this analysis is the endogenous adoption of laws. The areas that are friendly and less hostile to LGB workers may be the areas that are most likely to adopt sexual orientation anti-discrimination laws. The main threat to identification is that the timing of anti-discrimination laws is correlated with another factor causing an omitted variable bias. It could be that positive sentiment toward LGB workers causes both the law change and any change in labor market outcomes. We mitigate this by adopting a saturated triple-difference model, including all two-way fixed effects (county-year, year-SSP, and county-SSP), as proposed by (Burn, 2018). The inclusion of county-year fixed effects remove the need for county controls, addressing the lack of local support data for same-sex marriage. We refine the approach of (Burn, 2018) by employing the methodology of (Borusyak et al., 2021), designed for analyzing staggered policy adoptions like ours. This method accurately estimates the effects of anti-discrimination laws, considering their staggered implementation times.
@@
In Figure 2, we present the distribution of same-sex partnerships by state from 2005 to 2019, and in Table 2, we present the 10 counties with the largest share of same-sex partnerships. Unsurprisingly, the distribution of same-sex partnerships by state and by county is skewed toward more progressive states and counties with large cities in them that are known for having a large LGB population like San Francisco, the District of Columbia, New York City, and Boston. The LGB population varies from the heterosexual population in many dimensions including geography.
@@
We present descriptive statistics showing the differences in the same-sex couple population compared to the different-sex couple population broken down by education level and sex in Table 3. Descriptively, there is a clear difference in labor market outcomes and characteristics between men and women in same-sex partnerships and those in different-sex partnerships when controlling for education.
@@ 3 Identification Strategy & Estimation
$Y_{ipct}$ $=\alpha\;+\;\beta_{1}\bigl{[}\mathrm{SSP}_{ipct}\times\mathrm{ENDA}_{ct}\bigr{]}\;+\;\sum_{s=1}^{S}\lambda_{s}\bigl{[}\mathrm{SSP}_{ipct}\times\mathrm{SSM}_{st}\times\mathbf{1}_{\text{state}(c)=s}\bigr{]}$ (1)
$\quad+\theta_{ct}+\mu_{pc}+\kappa_{pt}+X_{ipct}^{\prime}\,\delta+\varepsilon_{ipct}.$
@@
In our regression, equation (1) defines the labor market outcomes $Y_{ipct}$ for individual $i$, in partnership type $p$, residing in county $c$, at time $t$. These outcomes include labor force participation, employment status, and the percentile rank of real wages (in 2019 dollars). The binary variable $\mathrm{SSP}_{ipct}$ indicates whether an individual is in a same‐sex partnership (1) or a different‐sex partnership (0). $\mathrm{SSM}_{st}$ indicates whether state $s$ recognizes same‐sex marriage in year $t$, and $\mathbf{1}_{\{\mathrm{state}(c)=s\}}=1$ if county $c$ lies in state $s$. The error term $\epsilon_{ipct}$ is clustered at the county level to account for within-county correlation. [^12] The ACS only identifies metro counties, so the county fixed effect is close to a pseudo-county fixed effect. Rural counties within a state are compared together with a true county fixed effect for metro area counties. Our key variable of interest is the interaction of same‐sex partnership status with the presence of a county‐level Employment Non‐Discrimination Act, $\mathrm{ENDA}_{ct}$. [^13] Treatment status is computed using the earliest year of ENDA implementation at the county or state level, with individuals remaining classified as “treated” thereafter.
We include three sets of high‐dimensional fixed effects to account for a wide range of unobserved factors. First, $\theta_{ct}$ (county×$\times$×year) absorbs any time‐varying shocks or policy changes common to all individuals within a county in a given year. Second, $\mu_{pc}$ (partnership×$\times$×county) controls for any unobserved, time‐invariant characteristics that are specific to each partnership type within a county. Third, $\kappa_{pt}$ (partnership×$\times$×year) accounts for year‐specific shocks that affect each partnership type differentially. The vector $X_{ipct}^{\prime}\delta$ includes additional demographic controls (e.g., race, age, and education). The coefficient of interest, $\beta_{1}$, identifies how ENDA coverage $\mathrm{ENDA}_{ct}$ differentially affects same‐sex partners, net of any confounding from same‐sex marriage legalization. The summation term $\sum_{s=1}^{S}\lambda_{s}\bigl{[}\mathrm{SSP}_{ipct}\times\mathrm{SSM}_{st}\bigr{]}$ partials out state‐specific labor‐market impacts of marriage legalization on same‐sex couples.
The key variation used to estimate $\beta_{1}$ is the differential timing of when counties adopted ENDAs, and how this interacted with the labor market outcomes for individuals in same-sex partnerships relative to those in different-sex partnerships. This variation allows us to isolate the effect of anti-discrimination laws on labor market outcomes by comparing same-sex and different-sex couples across counties with and without ENDAs, before and after their passage. Thus, $\beta_{1}$ captures the differential impact of ENDA on labor supply and wage outcomes specifically for individuals in same-sex partnerships. [^14] We recognize the heterogeneous impact (e.g., Burn (2018) of anti-discrimination laws across different regions, yet our analysis aims to estimate an average effect for those directly subject to the laws.
We limit our sample to people aged 25 to 65, and we estimate this model separately by sex using the ACS person weights. Following Chen and Roth (2024), we compute the percentile rank of annual and hourly real wages. [^15] Given the ACS’s 23-month reference period for earnings, as discussed by Hansen et al. (2020), this extended timeframe might not precisely align with the implementation dates of ENDAs, potentially biasing the immediate effects of these policies towards zero.